A 15-foot ladder is placed against the wall of a house such that the bottom of the ladder is at a distance of 10 feet from the wall.At what height does the ladder touch the house?
Answer by stanbon(75887)

Use Pythagoras. The hypotenuse is 15.
One leg is 10. Find the other leg.
15^2 = 10^2 + x^2
x^2 = 125
x = sqrt 125 = 5sqrt5
Cheers,
Stan H.